OpenStudy (anonymous):

Write the expression sin(tan^-1)x) as an algebraic expression in x (without trig or inverse trig functions). A. (sqrt (1+x^2))/ (x) B. (1) / (sqrt (1+x^2)) C. (x) / (sqrt (1-x^2)) D. (x) / (sqrt (1+x^2)) E. (1) / (sqrt (1-x^2))

OpenStudy (anonymous):

make a triangle... the ans will be obvious. it is opp/hyp so it is x/(x^2+1)

OpenStudy (anonymous):

Ah a triangle! Thank you!

OpenStudy (anonymous):

yes a triangle. label "opposite" side x and adjacent side 1 since the tangent is x = x/1 use pythatoras to find the hypotenuse which will be \[\sqrt{1+x^2}\]

OpenStudy (anonymous):

then \[sin(tan^{-1}(x))=\frac{x}{\sqrt{1+x^2}}\]

OpenStudy (anonymous):

that being opposite over hypotenuse

OpenStudy (anonymous):

and by the same triangle \[cos(tan^{-1}(x))=\frac{1}{\sqrt{1+x^2}}\]